The length of a rectangle is 8 in. more than its width. The perimeter of the rectangle is 24 in. What are the width and length of the rectangle?
Found 2 solutions by mananth, solver91311:
Answer by mananth(16946)

The length of a rectangle is 8 in. more than its width. The perimeter of the rectangle is 24 in. What are the width and length of the rectangle?
...
let width be x in
length = x+8 in.
perimeter = 2*(L+W)
2*(x+x+8)=24
2(2x+8)=24
4x+16=24
4x=24-16
4x=8
/4
x=2 in the width
x=2 = 2+8 =10 in the length
...
m.ananth@hotmail.ca

Answer by solver91311(24713)





Solve for

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism